OpenStudy (anonymous):

WHAT IS RADICAL 50 TO THE FIFTH POWER. (SIMPLIFIED)

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

@KingGeorge

OpenStudy (anonymous):

@robtobey

OpenStudy (lgbasallote):

im not even tagged? :(

OpenStudy (anonymous):

can you please help please :( ugh @lgbasallote

OpenStudy (lgbasallote):

do you mean \[\huge \sqrt[5]{50}\]or\[\huge \sqrt{(50)^5}\]

OpenStudy (anonymous):

the second one @lgbasallote

OpenStudy (lgbasallote):

first of all... \[\huge \begin{array} &\sqrt x^2 \implies x \\ \sqrt{x^3} \implies x\sqrt{x^2} \\ \sqrt{x^4} \implies x^2 \end{array}\] so what do you think is the value of \(\large \sqrt{x^5}\)

OpenStudy (anonymous):

i have no idea @lgbasallote

OpenStudy (anonymous):

is it the first one?